### v3

@@ 5 Algorithm Design
In this section, we present our algorithms for solving the CMDP problem. We first show how to identify one optimal basis of LP (8) and then present the formal algorithm to achieve a logarithmic level regret bound.
5.1 Characterize One Optimal Basis
We now describe how to identify one optimal basis of the LP (8), by sequentially discarding the sub-optimal actions and the redundant constraints. We will use the estimates constructed in the previous Section 4 to carry out the discarding procedure. The formal algorithm to identify such non-zero elements and the constraints is given in Algorithm 1.
@@ 5.1 Characterize One Optimal Basis
1:Input: the historical sample set $\mathcal{F}_{N_{0}}$ that contains $N_{0}$ samples for each $(s,a)\in\mathcal{S}\times\mathcal{A}$.
2:Compute the value of $\bar{V}_{N_{0}}$ as in (10).
3:Initialize $\mathcal{I}$ to be the whole index set that contains every column index of matrix $B$ in (8) and $\mathcal{J}=[K]$.
4:for $i\in\mathcal{I}$ do
5:     Let =\$\𝑖\mathcal{I}^{\prime}=\mathcal{I}\backslash\{i\}$.
@@
7:     If $|\bar{V}_{\mathcal{I}^{\prime},N_{0}}-\bar{V}_{N_{0}}|\leq 2\mathrm{Gap}_{1}(N_{0},\varepsilon)+2\mathrm{Gap}_{2}(N_{0},\varepsilon)$, then we set $\mathcal{I}=\mathcal{I}^{\prime}$.
@@
9:for $k=1,\dots,K$ do
10:     Let $\mathcal{J}^{\prime}=\mathcal{J}\backslash\{q\}$.
11:     Compute the value of $\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I},N_{0}}$ as in (20).
12:     If $|\bar{V}_{N_{0}}-\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I},N_{0}}|\leq 2\mathrm{Gap}_{1}(N_{0},\varepsilon)+2\mathrm{Gap}_{2}(N_{0},\varepsilon)$, then we set $\mathcal{J}=\mathcal{J}^{\prime}$.
13:end for
14:Output: the set of indexes $\mathcal{I}$ and $\mathcal{J}$.
We now explain the intuition why Algorithm 1 works. Denote by $\bm{q}^{*}$ an optimal solution to LP (8). We describe how to identify the non-zero elements in $\bm{q}^{*}$ and how to identify the constraints such that the values of the non-zero elements of $\bm{q}^{*}$ can be uniquely determined by the corresponding linear equation. For each $i$-th element of $\bm{q}^{*}$, we compare the value of $V$ (8) against $V$ with an additional constraint that $q_{i}=0$. If the two values are different, we identify a non-zero element. To this end, for an index set $\mathcal{I}$, we define an LP, as well as its estimate, as follows.
$V_{\mathcal{I}}=$ $\max$ $\hat{\bm{r}}^{\top}\bm{q}$ $\bar{V}_{\mathcal{I},N_{0}}=$ $\max$ $(\bar{\bm{r}}_{N_{0}})^{\top}\bm{q}$ (17)
    s.t. $C\bm{q}\leq\bm{\alpha}$     s.t. $\bar{C}_{N_{0}}\bm{q}\leq\bm{\alpha}+\lambda_{N_{0}}$
$B\bm{q}=\bm{\mu}$ $|\bar{B}_{N_{0}}\bm{q}-\bm{\mu}|\leq\lambda_{N_{0}}$
@@
where $\mathcal{I}^{c}$ denotes the complementary set of $\mathcal{I}$. Note that if $V-V_{\mathcal{I}}>0$, we know that $\mathcal{I}^{c}$ contains a non-zero basic variable. The steps 4-8 in Algorithm 1 reflect this point. Starting from $\mathcal{I}$ denoting the whole index set, we sequentially delete one element $i$ (denoting $(s,a)$ in the infinite horizon discounted problem) from the set $\mathcal{I}$. Once we detected that $V-V_{\mathcal{I}\backslash\{i\}}>0$, we know that $i$ is a non-zero basic variable and we add $i$ back into the set $\mathcal{I}$. In this way, we can classify all the basic variables into the set $\mathcal{I}$. However, since we do not know the exact value of $V$ and $V_{\mathcal{I}}$, we use the estimates and compare the value of $\bar{V}_{N_{0}}$ and $\bar{V}_{\mathcal{I},N_{0}}$. For this comparison to be valid, the estimation error has to be smaller than the intrinsic gap between $V$ and $V_{\mathcal{I}}$. We define a constant
@@
and we need $N_{0}$ to be large enough such that the estimation gap is smaller than $\delta_{1}/2$. The quantity $\delta_{1}$ plays a similar role as the gap between the optimal arm and the sub-optimal arm in the multi-arm-bandit problem when deriving a problem-dependent bound (e.g. Lai and Robbins (1985)).
Denote by $M$ the number of non-zero basic variables detected in Algorithm 1. We then denote by $M^{\prime}$ the number of binding constraints corresponding to the optimal solution given by the optimal basis we detected. Our goal is to identify the linearly independent binding constraints such that the non-zero basic variables can be uniquely determined as the solution to a linear equation, as described in Lemma 3.1.
To tell which constraints are redundant and thus can be deleted, we consider the dual program of $V_{\mathcal{I}}$, where $\mathcal{I}$ is determined in steps 4-8 in Algorithm 1, and similarly, we test which dual variable can be set to 00 without influencing the dual objective value. For a dual variable index subset $\mathcal{J}\subset[K]$, we consider the dual program as follows.
$\text{Dual}_{\mathcal{J},\mathcal{I}}=\min\leavevmode\nobreak\ \leavevmode\nobreak$ $\bm{\alpha}^{\top}\bm{y}+\bm{\mu}^{\top}\bm{z}$ (19a)
$\bm{y}_{\mathcal{J}^{c}}=0$ (19c)
$\bm{y}\geq 0,\bm{z}\geq-\infty,$ (19d)
where $\mathcal{J}^{c}=[K]\backslash\mathcal{J}$. The estimate of $\text{Dual}_{\mathcal{J},\mathcal{I}}$, can be obtained from the estimate of its dual, which is given below.
$\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}=\bar{V}_{\mathcal{J},\mathcal{I},N_{0}}=\max\leavevmode\nobreak\ \leavevmode\nobreak$ $(\bar{\bm{r}}_{N_{0}})^{\top}\bm{q}$ (20a)
s.t. $\bar{C}_{N_{0}}(\mathcal{J},:)\bm{q}\leq\bm{\alpha}+\lambda_{N_{0}}$ (20b)
$|\bar{B}_{N_{0}}\bm{q}-\bm{\mu}|\leq\lambda_{N_{0}}$ (20c)
$\bm{q}_{\mathcal{I}}=0$ (20d)
$\bm{q}\geq 0,$ (20e)
Similarly, we compare the value of $\text{Dual}_{\mathcal{I}}$, where $\text{Dual}_{\mathcal{I}}=\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I}}$ with $\mathcal{J}^{\prime}=[K]$, and $\text{Dual}_{\mathcal{J},\mathcal{I}}$. However, we can only compare the value of their estimates $\text{Dual}_{\mathcal{I},N_{0}}$ and $\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}$. To this end, we define a constant
$\delta_{2}=\min_{\mathcal{J}\subset[K]}\{\text{Dual}_{\mathcal{J},\mathcal{I}}-\text{Dual}_{\mathcal{I}}:\text{Dual}_{\mathcal{J},\mathcal{I}}-\text{Dual}_{\mathcal{I}}>0\}.$ (21)
For the comparison to be valid, we need $N_{0}$ to be large enough such that the estimation error is smaller than $\delta_{2}/2$. In this way, we identify the linearly independent binding constraints corresponding to $\bm{q}^{*}$, as described in steps 9-13 of Algorithm 1.

### v4

@@ 5 Algorithm Design
In this section, we present our algorithms for solving the CMDP problem. We first show how to identify one optimal basis of LP (8) and then present the formal algorithm to achieve an instance-dependent guarantee.
5.1 Optimal Basis Characterization (on True Problem)
To motivate our formal algorithm that identifies the optimal basis of the LP (8) from samples, we start with identifying one optimal basis assuming that we have the access to the true LP formulation in (8), which is equivalent to assuming that the number of samples $N_{0}\rightarrow\infty$ such that $\bar{C}_{N_{0}}=C$ and $\bar{B}_{N_{0}}=B$. The optimal basis of the LP (8) can actually be identified by sequentially discarding the sub-optimal actions and the redundant constraints. We present the full algorithm in Algorithm 1. We now explain the procedures in Algorithm 1 and show Algorithm 1 successfully identify one optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$.
One key design in our algorithm is to consider the LP $V_{\mathcal{I}}$ defined for a set $\mathcal{I}\subset\mathcal{S}\times\mathcal{A}$ as follows
[…]
s.t. $C\bm{q}\leq\bm{\alpha}$
$B\bm{q}=\bm{\mu}$
$\bm{q}_{\mathcal{I}^{c}}=0$
$\bm{q}\geq 0.$
Note that if it holds $V=V_{\mathcal{I}}$ for a set $\mathcal{I}$, then it implies that the set $\mathcal{I}$ contains a set of basic variables for the original LP $V$ such that restricting the elements of the decision variable $\bm{q}$ in the set $\mathcal{I}^{c}$ to be 0 will not change the LP value. Following this intuition, starting from $\mathcal{I}$ denoting the whole index set, we sequentially delete one element $i$ (denoting one state-action pair $(s,a)$ in the infinite horizon discounted problem) from the set $\mathcal{I}$. Once we detected that $V\{$$V_{\mathcal{I}\backslash\{i\}}>0$, we know that $i$ is a non-zero basic variable and we remain $i$ in the set $\mathcal{I}$. If we detected that $V\{$$V_{\mathcal{I}\backslash\{i\}}$, we know that we can safely drop the element $i$ from the set $\mathcal{I}$ and we will do that permanently, which implies that the dropped elements will never be added back to the set $\mathcal{I}$. In this way, we can classify all the non-zero basic variables into the set $\mathcal{I}$. This describes the procedure of step 3-8 in Algorithm 1.
Algorithm 1 Algorithm for identifying one optimal basis on the true problem
1:Input: the true LP $V$.
2:Set $\mathcal{I}_{0}$ to be the whole index set of $\mathcal{S}\times\mathcal{A}$ and set $\mathcal{J}_{0}=[K]\cup\mathcal{S}$.
3:Initialize $\mathcal{I}=\mathcal{I}_{0}$ and $\mathcal{J}=\mathcal{J}_{0}$.
4:for $i\in\mathcal{I}_{0}$ do
[…]
6:  Compute the value of $V_{\mathcal{I}^{\prime}}$ as in (19).
7:  If $V_{\mathcal{I}^{\prime}}=V$, then we set $\mathcal{I}=\mathcal{I}^{\prime}$.
8:end for
9:for $j\in\mathcal{J}_{0}$ do
[…]
11:  Compute the value of $\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I}}$ as in (18).
12:  If $V=\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I}}$ and the matrix $\begin{bmatrix}C(\mathcal{J}^{\prime}_{1},\mathcal{I})\\ B(\mathcal{J}^{\prime}_{2},\mathcal{I})\end{bmatrix}$ is full rank, then we set $\mathcal{J}=\mathcal{J}^{\prime}$.
13:  Break if $|\mathcal{J}|=|\mathcal{I}|$.
14:end for
15:Output: the set of indexes $\mathcal{I}$ and $\mathcal{J}$.
We now aim to obtain the supporting constraints. Denote by $M$ the number of non-zero basic variables detected in Algorithm 1. We then denote by $M^{\prime}$ the number of binding constraints corresponding to the optimal solution given by the optimal basis we detected. Our goal is to identify the linearly independent binding constraints such that the non-zero basic variables can be uniquely determined as the solution to $M$ number of linear equations, as described in Lemma 3.1. To tell which constraints are redundant and thus can be deleted, we consider the dual program of $V_{\mathcal{I}}$, where $\mathcal{I}$ is determined in steps 3-8 in Algorithm 1, and similarly, we test which dual variable can be set to 0 without influencing the dual objective value. For a dual variable index subset $\mathcal{J}=(\mathcal{J}_{1},\mathcal{J}_{2})$ where $\mathcal{J}_{1}\subset[K]$ and $\mathcal{J}_{2}\subset\mathcal{S}$, we consider the dual program as follows.
$\text{Dual}_{\mathcal{J},\mathcal{I}}=\min\penalty 10000\ \penalty 10000\$$\bm{\alpha}^{\top}\bm{y}+\bm{\mu}^{\top}\bm{z}$ (18a)
$\bm{y}_{\mathcal{J}_{1}^{c}}=0,\penalty 10000\ \bm{z}_{\mathcal{J}_{2}^{c}}=0,$ (18c)
$\bm{y}\geq 0,\bm{z}\geq-\infty,$ (18d)
where $\mathcal{J}_{1}^{c}=[K]\backslash\mathcal{J}_{1}$ and $\mathcal{J}_{2}^{c}=\mathcal{S}\backslash\mathcal{J}_{2}$.
[…]
Theorem 5.1
For the true LP $V$, the sets $\mathcal{I}$ and $\mathcal{J}$ output by Algorithm 1 satisfy the conditions in Lemma 3.1.
The proof of Theorem 5.1 is relegated to Section 12.1. We now in the following remark address the main differences between our Algorithm 1 and other algorithms in the previous literature.
Remark 5.2
[…]
Note that there are also other algorithms existing in the literature that can identify an optimal basis of an LP (e.g. megiddo1991finding). However, as illustrated in Remark 3.2, our Algorithm 1 is not to find exactly an optimal basis, instead, we find the set of non-zero basic variables and the corresponding supporting constraints. When the LP $V$ is degenrate, the basic variables founded by the algorithm in megiddo1991finding can be zero, which would cause troubles when doing the online learning extension in . Moreover, our Algorithm 1 only involves operations such as checking the LP values and the non-singularity of a matrix, which are easy to be extended to the data-driven setting where we only have sample average estimations of the constraint matrix. In general, our Algorithm 1 is new in the literature and brings advantages as discussed above to handling the problem in our settings.
5.2 Characterize One Optimal Basis From Samples
We now describe how to carry out our Algorithm 1 when we don’t have access to the true LP $V$. We will use the estimates constructed in the previous Section 4 to carry out the procedure. The formal algorithm is presented in Algorithm 2.
@@ 5.2 Characterize One Optimal Basis From Samples
1:Input: set $\mathcal{F}_{N_{0}}$ containing $N_{0}$ samples for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ and failure probability $\epsilon$.
2:Set $\mathcal{I}_{0}$ to be the whole index set of $\mathcal{S}\times\mathcal{A}$ and set $\mathcal{J}_{0}=[K]\cup\mathcal{S}$, and compute the value of $\bar{V}_{N_{0}}$ as in (10).
3:Initialize $\mathcal{I}=\mathcal{I}_{0}$ and $\mathcal{J}=\mathcal{J}_{0}$.
4:for $i\in\mathcal{I}_{0}$ do
[…]
@@
7:  If $|\bar{V}_{\mathcal{I}^{\prime},N_{0}}-\bar{V}_{N_{0}}|\leq 2\mathrm{Gap}_{1}(N_{0},\epsilon)+2\mathrm{Gap}_{2}(N_{0},\epsilon)$, then we set $\mathcal{I}=\mathcal{I}^{\prime}$.
@@
9:for $j\in\mathcal{J}_{0}$ do
[…]
11:  Compute $\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I},N_{0}}$ as in (19), and the smallest singular value $\bar{\sigma}$ of $\begin{bmatrix}\bar{C}_{N_{0}}(\mathcal{J}^{\prime}_{1},\mathcal{I})\\ \bar{B}_{N_{0}}(\mathcal{J}^{\prime}_{2},\mathcal{I})\end{bmatrix}$.
12:  if $|\bar{V}_{N_{0}}-\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I},N_{0}}|\leq 2\mathrm{Gap}_{1}(N_{0},\epsilon)+2\mathrm{Gap}_{2}(N_{0},\epsilon)$ and $\bar{\sigma}\geq|\mathcal{J}^{\prime}||\mathcal{I}|\cdot\mathrm{Rad}(N_{0},\epsilon)$ then
13:   set $\mathcal{J}=\mathcal{J}^{\prime}$.
14:  end if
15:  Break if $|\mathcal{J}|=|\mathcal{I}|$.
16:end for
17:Output: the set of indexes $\mathcal{I}$ and $\mathcal{J}$.
We now explain why Algorithm 2 works. The Algorithm 2 is developed based on Algorithm 1. From Theorem 5.1, we know that the Algorithm 1 would be able to find the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$. Therefore, as long as the Algorithm 2 follows the same procedure of Algorithm 1 by dropping exactly the same element $i$ from the set $\mathcal{I}$ and the same element $j$ from the set $\mathcal{J}$ as Algorithm 1, we know that Algorithm 2 finds the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$. However, the true LP $V$ is unknown to the Algorithm 2. Therefore, the Algorithm 2 focuses on the sample average estimation $\bar{V}_{\mathcal{I},N_{0}}$ of $V_{\mathcal{I}}$, given as follows, as well as the sample average estimation $\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}$ of $\text{Dual}_{\mathcal{J},\mathcal{I}}$, given as follows.
$\bar{V}_{\mathcal{I},N_{0}}=$ $\max$ $(\bar{\bm{r}}_{N_{0}})^{\top}\bm{q}$ $\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}=\bar{V}_{\mathcal{J},\mathcal{I},N_{0}}=$ $\max$ $(\bar{\bm{r}}_{N_{0}})^{\top}\bm{q}$ (19)
  s.t. $\bar{C}_{N_{0}}\bm{q}\leq\bm{\alpha}+\lambda_{N_{0}}$   s.t. $\bar{C}_{N_{0}}(\mathcal{J}_{1},:)\bm{q}\leq\bm{\alpha}_{\mathcal{J}_{1}}+\lambda_{N_{0}}$
$\|\bar{B}_{N_{0}}\bm{q}-\bm{\mu}\|_{\infty}\leq\lambda_{N_{0}}$ $\|\bar{B}_{N_{0}}(\mathcal{J}_{2},:)\bm{q}-\bm{\mu}\|_{\infty}\leq\lambda_{N_{0}}$
@@
where $\mathcal{I}^{c}$ denotes the complementary set of $\mathcal{I}$, and $\mathcal{J}_{1}=\mathcal{J}\cap[K]$, $\mathcal{J}_{2}=\mathcal{J}\cap\mathcal{S}$.
[…]
Therefore, we know that the condition in step 7 of Algorithm 2 that $|\bar{V}_{\mathcal{I}^{\prime},N_{0}}-\bar{V}_{N_{0}}|\leq 2\mathrm{Gap}_{1}(N_{0},\epsilon)+2\mathrm{Gap}_{2}(N_{0},\epsilon)$ is a necessary condition of $V=V_{\mathcal{I}^{\prime}}$ (with a high probability). In order to derive the sufficient condition of $V=V_{\mathcal{I}^{\prime}}$ (with a high probability), we need to introduce an intrinsic gap between $V$ and all possible $V_{\mathcal{I}^{\prime}}$, i.e., we define a constant
@@
It is clear to see that as long as the sample size $N_{0}$ is large enough such that $4\mathrm{Gap}_{1}(N_{0},\epsilon)+4\mathrm{Gap}_{2}(N_{0},\epsilon)<\delta_{1}$, the condition $|\bar{V}_{\mathcal{I}^{\prime},N_{0}}-\bar{V}_{N_{0}}|\leq 2\mathrm{Gap}_{1}(N_{0},\epsilon)+2\mathrm{Gap}_{2}(N_{0},\epsilon)$ would imply that $|V-V_{\mathcal{I}^{\prime}}|<\delta_{1}$. Given the definition of $\delta_{1}$ in (20), we must have $V=V_{\mathcal{I}^{\prime}}$. Therefore, we know that with a high probability, the step 7 of Algorithm 2 is exactly the same as the step 7 of Algorithm 1 and each element $i\in\mathcal{I}_{0}$ will be dropped/remained in a same way between Algorithm 2 and Algorithm 1.
To tell which constraints are redundant and thus can be deleted, we consider the estimate of the dual LP of $V_{\mathcal{I}}$, which is given as $\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}$ in (19). Then, the step 9-15 of Algorithm 2 carries out the same procedure as the step 9-14 of Algorithm 1. To see this point, we define a constant
$\delta_{2}=\min_{\mathcal{I},\mathcal{J}}\{\text{Dual}_{\mathcal{J},\mathcal{I}}-V_{\mathcal{I}}:\text{Dual}_{\mathcal{J},\mathcal{I}}-V_{\mathcal{I}}>0\}.$ (21)
Following the same procedure as above, it is easy to see that when the sample size $N_{0}$ is large enough such that $4\mathrm{Gap}_{1}(N_{0},\epsilon)+4\mathrm{Gap}_{2}(N_{0},\epsilon)<\delta_{2}$, the condition of $|\bar{V}_{N_{0}}-\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I},N_{0}}|\leq 2\mathrm{Gap}_{1}(N_{0},\epsilon)+2\mathrm{Gap}_{2}(N_{0},\epsilon)$ is a sufficient and necessary condition for $V=\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I}}$, with a high probability.
Moreover, we can apply the sensitivity of matrix singular values (e.g. Theorem 1 of stewart1998perturbation) to show that the gap between the smallest singular value of the matrix $\begin{bmatrix}\bar{C}_{N_{0}}(\mathcal{J}^{\prime}_{1},\mathcal{I})\\ \bar{B}_{N_{0}}(\mathcal{J}^{\prime}_{2},\mathcal{I})\end{bmatrix}$ and the matrix $\begin{bmatrix}C(\mathcal{J}^{\prime}_{1},\mathcal{I})\\ B(\mathcal{J}^{\prime}_{2},\mathcal{I})\end{bmatrix}$ is upper bounded by $|\mathcal{J}^{\prime}||\mathcal{I}|\cdot\mathrm{Rad}(N_{0},\epsilon)$, with a high probability. Therefore, denote by $\sigma_{0}$ the smallest possible singular value of the matrix $\begin{bmatrix}C(\mathcal{J}^{\prime}_{1},\mathcal{I})\\ B(\mathcal{J}^{\prime}_{2},\mathcal{I})\end{bmatrix}$ for all possible $\mathcal{J}^{\prime},\mathcal{I}$ such that the matrix $\begin{bmatrix}C(\mathcal{J}^{\prime}_{1},\mathcal{I})\\ B(\mathcal{J}^{\prime}_{2},\mathcal{I})\end{bmatrix}$ is full rank, i.e., we define the following constant
$\sigma_{0}=\min_{\mathcal{I},\mathcal{J}}\left\{\sigma_{\min}\left(\begin{bmatrix}C(\mathcal{J}_{1},\mathcal{I})\\ B(\mathcal{J}_{2},\mathcal{I})\end{bmatrix}\right)\text{\penalty 10000\ with\penalty 10000\ }\mathcal{J}_{1}=\mathcal{J}\cap[K],\mathcal{J}_{2}=\mathcal{J}\cap\mathcal{S}:\sigma_{\min}\left(\begin{bmatrix}C(\mathcal{J}_{1},\mathcal{I})\\ B(\mathcal{J}_{2},\mathcal{I})\end{bmatrix}\right)>0\right\}$ (22)
where $\sigma_{\min}(A)$ denotes the smallest singular value of a matrix $A$. We can show that as long as the sample size $N_{0}$ is large enough such that it holds $\sigma_{0}\geq 2|\mathcal{J}^{\prime}||\mathcal{I}|\cdot\mathrm{Rad}(N_{0},\epsilon)$, the condition that the smallest singular value of the matrix $\begin{bmatrix}\bar{C}_{N_{0}}(\mathcal{J}^{\prime}_{1},\mathcal{I})\\ \bar{B}_{N_{0}}(\mathcal{J}^{\prime}_{2},\mathcal{I})\end{bmatrix}$ to be lower bounded by $|\mathcal{J}^{\prime}||\mathcal{I}|\cdot\mathrm{Rad}(N_{0},\epsilon)$ is a sufficient and necessary condition for the matrix $\begin{bmatrix}C(\mathcal{J}^{\prime}_{1},\mathcal{I})\\ B(\mathcal{J}^{\prime}_{2},\mathcal{I})\end{bmatrix}$ to be full rank, with a high probability.
In this way, we know that with a high probability, the step 12 of Algorithm 2 is exactly the same as the step 12 of Algorithm 1 and each element $j\in\mathcal{J}_{0}$ will be dropped/remained in a same way between Algorithm 2 and Algorithm 1.